# Dropping Just a Handful of Preferences Can Change Top Large Language Model Rankings

### v2

@@ 1 Introduction
In this work, we study a different type of untrustworthiness of LLM ranking systems. That is: “Will the top rankings from LLM-evaluation platforms change upon dropping a very small fraction of the human (or AI) preference evaluations?” A positive answer would raise concerns about the stability and generalizability of rankings produced by such systems. Our notion of non-robustness differs from those of Min et al. (2025); Huang et al. (2025b); Zhao et al. (2025) in two major respects. First, it occurs at a different place in the process, at the data analysis step after data has been collected (including from malicious or apathetic users). Second, it does not require adversarial intent. Our notion is more concerned with statistical robustness, namely of a ranking learned from data to dropping a small fraction of the data. While we do aim to find a worst-case fraction, the intent is to provide an upper bound on the degree of non-robustness.
Our question posed above motivates the need for a systematic way to assess the robustness of top rankings in BT-based evaluation systems to worst-case data dropping. However, no such method currently exists, beyond a brute-force combinatorial search over all possible small subsets of data. [^1] This combinatorial search is computationally infeasible for large-scale platforms like Chatbot Arena. In order to avoid this computationally intractable search, we turn to a recent line of works from statistics and theoretical computer science that design algorithms for assessing whether data analyses are robust to dropping a small, worst-case fraction of data points (Broderick et al., 2020; Kuschnig et al., 2021; Moitra and Rohatgi, 2023; Freund and Hopkins, 2023; Shiffman et al., 2023; Nguyen et al., 2024; Huang et al., 2025a; Rubinstein and Hopkins, 2025). One such method, the Approximate Maximum Influence Perturbation (AMIP), estimates how much a statistic of interest could change if a worst-case subset of the data were dropped (Broderick et al., 2020). We extend these ideas to develop a fast approximation method for assessing the robustness of rankings from LLM evaluation systems to worst-case data-dropping.
We apply our method to assess several popular LLM ranking platforms, including Chatbot Arena and derivatives (Chiang et al., 2024b; Zheng et al., 2023; Miroyan et al., 2025; Vichare et al., 2025; Chou et al., 2025) and find most to be non-robust to dropping a very small fraction of votes.
@@ 2 Setup
$\hat{\bm{\theta}}$= $\operatorname*{arg\,max}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}\big[w_{WL}I_{y_{n}=W}\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+w_{WL}I_{y_{n}=L}\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))$ (3)
$\qquad\qquad\quad+w_{T}I_{y_{n}=T}\big(\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))\big)\big].$
@@
Finally, we define the rank of a model as its position in the sorted list of models, $(\widehat{\theta}_{(1)},\ldots,\widehat{\theta}_{(M)})$, ordered by their scores in descending order, so that $\widehat{\theta}_{(1)}$ corresponds to the top-ranked model.
@@
$\hat{\bm{\theta}}(w)$= $\operatorname*{arg\,max}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}w_{n}\big[w_{WL}I_{y_{n}=W}\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+w_{WL}I_{y_{n}=L}\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))$ (4)
$\quad\quad\qquad\qquad\quad+w_{T}I_{y_{n}=T}\big(\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))\big)\big].$
@@
We begin by analyzing the robustness of the ordering of BT scores between a pair of players, $i$ and $j$. Without loss of generality, we assume [^3] If this assumption does not hold, the identities of $i$ and $j$ can be swapped. that player $i$ has the higher estimated BT score on the full data:
@@
We now extend this notion to an arena with $M$ players, for any $M\geq 2$. Let $\mathcal{T}(w):=\{\widehat{\theta}_{i}(w)\}_{i=1}^{M}$ denote the set of BT scores under weighting $w$.
@@
The top-$k$ set under full data is defined as the set of players whose scores rank among the top $k$:
$\mathcal{K}_{\mathcal{T}(1_{N})}:=\left\{\widehat{\theta}_{i}(1_{N}):\operatorname{rank}\left[\widehat{\theta}_{i}(1_{N});\mathcal{T}(1_{N})\right]\leq k\right\}.$ (5)
@@
An arena is top-$k$ robust at level $\alpha$ if no $\alpha$-fraction subset of data can be dropped to change the top-$k$ set. That is,
@@
Notice that Equation 6 is nontrivial to directly verify; to check directly, we have to test out dropping all possible small-fraction subsets of the arena, a combinatorial operation that is computationally intractable in practice.
In Section 3, we show that verifying whether Equation 6 holds can be reduced to checking the robustness of a series of pairwise comparisons. Specifically, top-$k$ robustness as defined in Definition 3 can be checked by checking whether there exists a reweighting $w\in W_{\alpha}$ that flips the ranking of a pair $(i,j)$ such that $i$ is inside and $j$ is outside the top-$k$ set. We then can test if such flipping can happen by using a continuous approximation of the discrete weights $w$ (also known as “approximate data-dropping”) to identify a promising candidate subset of influential preferences, dropping these, recomputing the BT-based rankings, and observing whether the rankings change. We detail this procedure in Section 3.
@@ 3 Proposed method
Recall that our goal is to evaluate the robustness of the rankings induced by a BT-model $\{\widehat{\theta}_{(1)},...,\widehat{\theta}_{(M)}\}$ when a small fraction of matches (e.g., evaluations) is removed from the arena. To this end, we introduce a method based on checking the robustness of pairwise BT score differences. We provide pseudocode for our method in Algorithm 1 and explain its steps below.
@@
To evaluate the robustness of the sign of $\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})]$ to dropping a small fraction of matches, we adopt a recently-developed method from the statistics literature known as the Approximate Maximum Influence Perturbation (Broderick et al., 2020) (see Section C.3 for a more detailed discussion on how we adapt this method to our problem setup). This method approximates the maximal directional change in a statistic, e.g., $\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})]$, that can result from dropping a worst-case subset of data points (in our case, evaluations) of size at most $\left\lfloor\alpha N\right\rfloor$. This method allows us to sidestep running an expensive combinatorial search over all data subsets for the worst-case subset of matches to drop, a procedure that is computationally prohibitive for large LLM evaluation platforms like Chatbot Arena.
@@
We approximate this discrete optimization problem using AMIP approximation (Broderick et al., 2020), the idea is that, instead of solving the optimization directly, we first approximate the effect of dropping data by a first order Taylor expansion of the quantity $\widehat{\theta}_{i}(w)-\widehat{\theta}_{j}(w)$ over data weights $w$ and then solve the approximated optimization problem. In Appendix C, we provided a review of the general AMIP approximation, then formulate the both weighted and unweighted BT model as logistic regressions and explicit form of the approximation for BT models.
For a candidate pair of players, $(i,j)$, we check whether after dropping, $\widehat{\theta}_{i}(\tilde{w})-\widehat{\theta}_{j}(\tilde{w})]<0$. In other words, we refit the BT-model upon leaving out the subset of impactful evaluations identified by AMIP and check whether leaving out this subset induces a sign change in the difference of BT scores for the pair, $(i,j)$. We say that the BT scores for a pair of players, $(i,j)$, are non-robust if the sign of the difference in scores becomes negative upon refitting under $\tilde{w}$, (i.e., if $\widehat{\theta}_{i}(\tilde{w})-\widehat{\theta}_{j}(\tilde{w})]<0$).
@@
Thus, we check top-$k$ robustness by iterating over pairs of players. Note that, when checking the robustness of a given pair $(i,j)$, we allow matches between any two models (not only $(i,j)$) to be dropped. Since we only need to find one non-robust pair to render the set non-robust, not all pairs need to be checked. To save on compute, we take a greedy approach and start with comparing the most closely-ranked pairs between the top-$k$ ranked players and the remaining MkM-k players, where “closeness” is quantified using the absolute difference in BT scores fit on the full data. [^4] The robustness of the relative ranking of two players is correlated with the proximity of their BT scores as seen in Figure 18.; pairs with smaller BT-score gaps are more likely to exhibit data-dropping non-robustness. Upon finding any single pair that is pairwise non-robust at an $\alpha$-level, the procedure terminates early and returns the corresponding players and the indices of the dropped evaluations. We say that an arena is $\alpha$-level top-$k$ robust if there does not exist a pair of players $(i,j)$, where $i\in\mathcal{K}_{\mathcal{T}(w)}$ and $j\notin\mathcal{K}_{\mathcal{T}(w)}$, that are $\alpha$-level pairwise non-robust. While our method uses an approximation to identify the influential preferences, it then performs an exact recomputation of the Bradley–Terry scores with the identified preferences removed. As a result, all non-robustness reported in this paper is definitive: when we state that dropping $100\alpha\%$ of preferences changes the ranking, we have explicitly verified that the ranking does in fact change upon removal of the surfaced subset. However, the algorithm may not catch all cases of non-robustness (i.e., false negatives are possible). See Appendix H for an extended discussion on the possibility of false negatives.
@@
1:Input: Dataset $(X,y)$, rank $k$, drop fraction $\alpha$.
2:Output: (1) A determination of whether top-$k$ non-robustness was found. (2) If top-$k$ non-robustness is found, we additionally return the pair of players ($i,j$) whose rankings flipped, the differences in their scores pre- and post- data-dropping, and the most influential set $D_{\alpha}$.
@@
25:  Choose the $\left\lfloor\alpha N\right\rfloor$ values of $\Delta_{n}(i,j)$ that are the largest in the negative direction, assuming that $\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})>0$.
26:  $D_{\alpha}\leftarrow$ indices corresponding to the $\left\lfloor\alpha N\right\rfloor$ most negative $\Delta_{n}$ values.
@@
29:  $(\hat{\theta}_{i}(w)-\hat{\theta}_{j}(w))_{\text{AMIP}}\leftarrow(\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N}))+\sum_{n\in D_{\alpha}}\Delta_{n}$
@@
32:  $\hat{\bm{\theta}}(w)\leftarrow\text{FitBTModel}(X\setminus D_{\alpha},\,y\setminus D_{\alpha})$
@@
36:   return “Arena is $\alpha$-level top-$k$ non-robust”, $(i,j),(\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})),(\widehat{\theta}_{i}(w)-\widehat{\theta}_{j}(w)),D_{\alpha}$
@@
The above procedure is fast for assessing the robustness of preference-based ranking systems. For example, we tested our method on historical preference datasets released by the Chatbot Arena project and hosted on Hugging Face (Chiang et al., 2024b). Specifically, we run top-1 and top-5 robustness on a dataset of size around 50,00050{,}000 evaluations in under $3$ minutes on a personal computer equipped with an Apple M1 Pro CPU at 3200 MHz and 16 GB of RAM.
@@ 4.1 Data and Setup
We run our robustness check on a variety of LLM Arenas, including Chatbot Arena (Chiang et al., 2024b), MT-bench (Zheng et al., 2023), Search Arena (Miroyan et al., 2025), Webdev Arena (Vichare et al., 2025), and Vision Arena (Chou et al., 2025). For more information about each arena, see Appendix D. Our analysis relies on historical preference datasets released by the Chatbot Arena project (Chiang et al., 2024b) and publicly hosted on LMArena’s HuggingFace account. Each record represents a matchup consisting of two LLMs that answer the same prompt, the names of the two models, and the user label indicating preference for model A, model B, or a tie. Figure 3 presents the Bradley–Terry scores of the top-10 models on Chatbot Arena.
@@ 5.1 Vulnerabilities in AI Leaderboards
Despite its ease-of-use and widespread popularity, largescale, community-driven platforms like Chatbot Arena are found to be vulnerable to adversarial attacks that can distort model rankings. Min et al. (2025) demonstrate that Chatbot Arena is vulnerable to vote-rigging: by injecting just a few hundred manipulated votes (out of 1.7 million), attackers can significantly change the top model rankings. Similarly, Huang et al. (2025b) find that an attacker can accurately identify which model produced a response on Chatbot Arena, and use that to systematically upvote or downvote a target model and propose several defenses (e.g., authentication, rate limits, malicious-vote detection) that make the leaderboard more robust to adversarial agents. Injected votes may be especially easy to construct on LLM-as-a-judge systems, as recent works show that LLM judges can be gamed in systematic ways (Zheng et al., 2025; Raina et al., 2024). Beyond vote-rigging, Singh et al. (2025) identify other issues such as data leakage and private testing practices that allow large, proprietary model developers to selectively report the best-performing versions of their models on the arena. Zhao et al. (2025) present a case study showing that model rankings can shift when a fraction of votes comes from apathetic or arbitrary annotators. Their analysis finds that replacing $10\%$ of votes with uniform $\{0,1\}$ labels can move two models by up to five ranks. In contrast, we do not alter votes but instead demonstrate that rankings can change by removing an alarmingly small fraction ($0.0003\%$) of the votes. More importantly, while Zhao et al. (2025) present a case study focused on the rankings of three specific test models, we develop a systematic method to evaluate the robustness of BT-based ranking systems under worst-case data dropping, which also identifies the specific prompt–response pairs driving ranking flips. Finally, while all works in this section focus on Chatbot Arena, we extend our analysis to other domains (vision, web design, search, and multi-turn dialogue) and find the leaderboard rankings on these platforms to be similarly non-robust.

### v3

@@ 1 Introduction
In this work, we study a different type of untrustworthiness of LLM ranking systems. That is: “Will the top rankings from LLM-evaluation platforms change upon dropping a very small fraction of the human (or AI) preference evaluations?” A positive answer would raise concerns about the stability and generalizability of rankings produced by such systems. Our notion of non-robustness differs from those of Min et al. (2025); Huang et al. (2025b); Zhao et al. (2025) in two major respects. First, it occurs at a different place in the process, at the data analysis step after data has been collected (including from malicious or apathetic users). Second, it does not require adversarial intent.
Our question posed above motivates the need for a systematic way to assess the robustness of top rankings in BT-based evaluation systems to worst-case data dropping. However, a brute-force combinatorial search over all possible small subsets of data would be computationally infeasible for large-scale platforms like Chatbot Arena. So we instead turn to a recent line of works from statistics and theoretical computer science that design algorithms for assessing whether data analyses are robust to dropping a small, worst-case fraction of data points (Broderick et al., 2020; Kuschnig et al., 2021; Moitra and Rohatgi, 2023; Freund and Hopkins, 2023; Shiffman et al., 2023; Nguyen et al., 2024; Huang et al., 2025a; Rubinstein and Hopkins, 2025). One such method, the Approximate Maximum Influence Perturbation (AMIP), estimates how much a statistic of interest could change if a worst-case subset of the data were dropped (Broderick et al., 2020). We extend these ideas to develop a fast approximation method for assessing the robustness of rankings from LLM evaluation systems to worst-case data-dropping.
We apply our method to assess several popular LLM ranking platforms, including Chatbot Arena and derivatives (Chiang et al., 2024b; Zheng et al., 2023; Miroyan et al., 2026; Vichare et al., 2025; Chou et al., 2025) and find most to be non-robust to dropping a very small fraction of votes.
@@ 2 Setup
Finally, the rank of a model is its position in the sorted list of models, $(\widehat{\theta}_{(1)},\ldots,\widehat{\theta}_{(M)})$, ordered by their scores in descending order, so that $\widehat{\theta}_{(1)}$ corresponds to the top-scoring and top-ranked model. [^1] For ease of exposition in the main text, we assume there are no ties. In practice, we did not observe ties in Bradley–Terry-based point estimates. For rankings based on confidence intervals, we do observe and handle ties; see Section A.1 for details.
@@
$\hat{\bm{\theta}}$:= $\operatorname*{arg\,max}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}\big[w_{WL}I_{y_{n}=W}\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+w_{WL}I_{y_{n}=L}\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))$
$\qquad\qquad\quad+w_{T}I_{y_{n}=T}\left\{\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))\right\}\big].$ (3)
@@
$\hat{\bm{\theta}}(w):=$ $\operatorname*{arg\,max}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}w_{n}\big[w_{WL}I_{y_{n}=W}\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+w_{WL}I_{y_{n}=L}\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))$
$\quad\quad\qquad\qquad\quad+w_{T}I_{y_{n}=T}\left\{\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}})\big)\right\}\big].$ (4)
@@
We begin by considering the ordering of BT scores between a pair of players, $i$ and $j$. Without loss of generality, we assume [^3] If this assumption does not hold, the identities of $i$ and $j$ can be swapped. that player $i$ has the higher estimated BT score on the full data:
@@
We now extend this notion to an arena with $M$ players, for any $M\geq 2$. Let $\mathcal{T}(w):=\{\widehat{\theta}_{i}(w)\}_{i=1}^{M}$ denote the set of BT scores under weighting $w$. Let $\operatorname{rank}\left[\widehat{\theta}_{i}(w);\mathcal{T}(w)\right]$ denote the rank of a model under the weighting $w$.
@@
We define the top-$k$ set under a data weighting $w$ as the set of players whose scores rank among the top $k$:
$\mathcal{K}_{\mathcal{T}(w)}:=\left\{i:\operatorname{rank}\left[\widehat{\theta}_{i}(w);\mathcal{T}(w)\right]\leq k\right\}.$ (5)
@@
An arena is top-$k$ robust at level $\alpha$ if no $\alpha$-fraction subset of data can be dropped to change the top-$k$ set for the full data. That is,
@@
Notice that Equation 6 is nontrivial to directly verify; to check directly, we could test out dropping all possible small-fraction subsets of the arena, but this combinatorial operation is computationally intractable in practice.
In Section 3, we show that verifying whether Equation 6 holds can be reduced to checking the robustness of a series of pairwise comparisons. Specifically, top-$k$ robustness as in Definition 3 can be checked by assessing whether there exists a reweighting $w\in W_{\alpha}$ that flips the ranking of a pair $(i,j)$ such that $i$ is inside and $j$ is outside the top-$k$ set. We then can test if such flipping can happen by: (a) using a continuous approximation of the discrete weights $w$ (also known as “approximate data-dropping”) to identify a promising candidate subset of influential preferences, (b) dropping these, (c) recomputing the BT-based rankings, and (d) observing whether the rankings change. We detail this procedure in Section 3.
@@ 3 Proposed method
Recall that our goal is to evaluate the robustness of the rankings induced by a BT-model when a small fraction of matches (e.g., evaluations) is removed from the arena. To this end, we introduce a method based on checking the robustness of pairwise BT score differences. We provide pseudocode for our method in Algorithm 1 and explain its steps below.
@@
To evaluate the robustness of the sign of $\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})]$ to dropping a small fraction of matches, we adopt a recently-developed method from the statistics literature known as the Approximate Maximum Influence Perturbation (Broderick et al., 2020) (see Section C.3 for a more detailed discussion on how we adapt this method to our problem setup). This method approximates the maximal directional change in a statistic, e.g., $\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})]$, that can result from dropping a worst-case subset of data points (in our case, evaluations) of size at most $\lfloor\alpha N\rfloor$. This method allows us to sidestep running an expensive combinatorial search over all data subsets for the worst-case subset of matches to drop, a procedure that is computationally prohibitive for large LLM evaluation platforms like Chatbot Arena.
@@
We approximate this discrete optimization problem using AMIP approximation (Broderick et al., 2020); the idea is that, instead of solving the optimization directly, we first approximate the effect of dropping data by a first order Taylor expansion of the quantity $\widehat{\theta}_{i}(w)-\widehat{\theta}_{j}(w)$ over data weights $w$ and then solve the approximated optimization problem. In Appendix C, we provide a review of the general AMIP approximation, then formulate both the weighted and unweighted BT models as logistic regressions, and finally provide an explicit form of the approximation for both BT models.
Let the approximate solution to Equation 8 that is returned by AMIP be denoted as $\tilde{w}$ (i.e., the set of data weights that are 0 at indices of data points that AMIP chooses to drop and 1 elsewhere). For a candidate pair of players, $(i,j)$, we check whether after dropping, $\widehat{\theta}_{i}(\tilde{w})-\widehat{\theta}_{j}(\tilde{w})]<0$. In other words, we refit the BT-model upon leaving out the subset of impactful evaluations identified by AMIP and check whether leaving out this subset induces a sign change in the difference of BT scores for the pair, $(i,j)$. We say that the BT scores for a pair of players, $(i,j)$, are non-robust if the sign of the difference in scores becomes negative upon refitting under $\tilde{w}$, (i.e., if $\widehat{\theta}_{i}(\tilde{w})-\widehat{\theta}_{j}(\tilde{w})]<0$).
@@
Thus, we check top-$k$ robustness by iterating over pairs of players. Note that, when checking the robustness of a given pair $(i,j)$, we allow matches between any two models (not only $(i,j)$) to be dropped. Since we only need to find one non-robust pair to render the set non-robust, not all pairs need to be checked. To save on compute, we take a greedy approach and start with comparing the most closely-ranked pairs between the top-$k$ ranked players and the remaining MkM-k players, where “closeness” is quantified using the absolute difference in BT scores fit on the full data; [^4] The robustness of the relative ranking of two players is correlated with the proximity of their BT scores as seen in Figure 18. pairs with smaller BT-score gaps are more likely to exhibit data-dropping non-robustness. Upon finding any single pair that is pairwise non-robust at an $\alpha$-level, the procedure terminates early and returns the corresponding players and the indices of the dropped evaluations. We say that an arena is $\alpha$-level top-$k$ robust if there does not exist a pair of players $(i,j)$, where $i\in\mathcal{K}_{\mathcal{T}(w)}$ and $j\notin\mathcal{K}_{\mathcal{T}(w)}$, that are $\alpha$-level pairwise non-robust. While our method uses an approximation to identify the influential preferences; it then performs an exact recomputation of the Bradley–Terry scores with the identified preferences removed. As a result, all non-robustness reported in this paper is definitive: when we state that dropping $100\alpha\%$ of preferences changes the ranking, we have explicitly verified that the ranking does in fact change upon removal of the surfaced subset. However, the algorithm may not catch all cases of non-robustness (i.e., false negatives are possible). See Appendix H for an extended discussion on the possibility of false negatives.
@@
1:Input: Dataset $\mathcal{D}$ (the collection of matches, e.g., preferences), rank $k$, drop fraction $\alpha$.
2:Output: (1) A determination of whether top-$k$ non-robustness was found. (2) If top-$k$ non-robustness is found, we additionally return the pair of players ($i,j$) whose rankings flipped, the differences in their scores pre- and post- data-dropping, and the most influential set II.
@@
25:  Choose the $\lfloor\alpha N\rfloor$ values of $\Delta_{n}(i,j)$ that are the largest in the negative direction, assuming that $\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})>0$.
26:  $I\leftarrow$ indices corresponding to the $\lfloor\alpha N\rfloor$ most negative $\Delta_{n}$ values.
@@
29:  $(\hat{\theta}_{i}(w)-\hat{\theta}_{j}(w))_{\text{AMIP}}\leftarrow(\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N}))+\sum_{n\in I}\Delta_{n}$
@@
32:  $\hat{\bm{\theta}}(w)\leftarrow\text{FitBTModel}(\mathcal{D}_{-\alpha})$, where $\mathcal{D}_{-\alpha}$ is the set of points in $\mathcal{D}$ except those with indices in II
@@
36:   return “Arena is $\alpha$-level top-$k$ non-robust”, $(i,j),(\widehat{\theta}_{i}(1_{N})-\widehat{\theta}_{j}(1_{N})),(\widehat{\theta}_{i}(w)-\widehat{\theta}_{j}(w)),I$
@@
The above procedure is fast for assessing the robustness of preference-based ranking systems. For example, we tested our method on historical preference datasets released by the Chatbot Arena project and hosted on Hugging Face (Chiang et al., 2024b). Specifically, we check top-1 and top-5 robustness on a dataset of size around 50,00050{,}000 evaluations in under $3$ minutes on a personal computer equipped with an Apple M1 Pro CPU at 3200 MHz and 16 GB of RAM.
@@ 4.1 Data and Setup
We run our robustness check on a variety of LLM Arenas, including Chatbot Arena (Chiang et al., 2024b), MT-bench (Zheng et al., 2023), Search Arena (Miroyan et al., 2026), Webdev Arena (Vichare et al., 2025), and Vision Arena (Chou et al., 2025). For more information about each arena, see Appendix D. Our analysis relies on historical preference datasets released by the Chatbot Arena project (Chiang et al., 2024b) and publicly hosted on LMArena’s HuggingFace account. Each record represents a matchup consisting of two LLMs that answer the same prompt, the names of the two models, and the user label indicating preference for model A, model B, or a tie. Figure 3 presents the Bradley–Terry scores of the top-10 models on Chatbot Arena.
@@ 5.1 Vulnerabilities in AI Leaderboards
Despite its ease-of-use and widespread popularity, large-scale, community-driven platforms like Chatbot Arena are found to be vulnerable to adversarial attacks that can distort model rankings. Min et al. (2025) demonstrate that Chatbot Arena is vulnerable to vote-rigging: by injecting just a few hundred manipulated votes (out of 1.7 million), attackers can significantly change the top model rankings. Similarly, Huang et al. (2025b) find that an attacker can accurately identify which model produced a response on Chatbot Arena, and use that to systematically upvote or downvote a target model. They propose several defenses (e.g., authentication, rate limits, malicious-vote detection) that make the leaderboard more robust to adversarial agents. Injected votes may be especially easy to construct on LLM-as-a-judge systems, as recent works show that LLM judges can be gamed in systematic ways (Zheng et al., 2025; Raina et al., 2024). Beyond vote-rigging, Singh et al. (2025) identify other issues such as data leakage and private testing practices that allow large, proprietary model developers to selectively report the best-performing versions of their models on the arena. Zhao et al. (2025) present a case study showing that model rankings can shift when a fraction of votes comes from apathetic or arbitrary annotators. Their analysis finds that replacing $10\%$ of votes with uniform $\{0,1\}$ labels can move two models by up to five ranks. In contrast, we do not alter votes but instead demonstrate that rankings can change by removing an alarmingly small fraction ($0.003\%$) of the votes. More importantly, while Zhao et al. (2025) present a case study focused on the rankings of three specific test models, we develop a systematic method to evaluate the robustness of BT-based ranking systems under worst-case data dropping, which also identifies the specific prompt–response pairs driving ranking flips. Beyond pairwise-preference-based rankings, past works have pointed out the fragility of LLM ranking systems based on absolute performance scores. Perlitz et al. (2024) show that rankings derived from the Mean Win Rate (MWR) metric (a variant of the Borda count) can be gamed to increase the ranking of a target model by evaluating numerous models that are almost equal to, but slightly weaker than, the target model, thereby inflating the target model’s ranking. The authors suggest allocating more evaluation resources for models ranked at the top of the leaderboard, where the rankings may carry more weight. Finally, while all pairwise-preference-based related works in this section focus on Chatbot Arena, we extend our analysis to other domains (vision, web design, search, and multi-turn dialogue) and find the leaderboard rankings on these platforms to be similarly non-robust.